At the end of the year, a company offered to buy 4,260 units of a product from X Company for a special price of $12.00 each instead of the company's regular price of $19.00 each. The following information relates to the 65,500 units of the product that X Company made and sold to its regular customers during the year:

Per-Unit Total     
Cost of goods sold $9.38    $614,390   
Period costs 2.70    176,850   
Total $12.08    $791,240   


Fixed cost of goods sold for the year were $140,170, and fixed period costs were $81,875. Variable period costs include selling commissions equal to 4% of revenue.

6. Profit on the special order is

Incorrect. Tries 3/3 Previous Tries


7. Assume the following two changes for the special order: 1) variable cost of goods sold will decrease by $0.80 per unit, and 2) there will be no selling commissions. What would be the effect of these two changes on the special order profit?

Tries 0/3


8. There is concern that regular customers will find out about the special order, and X Company's regular sales will fall by 700 units. As a result of these lost sales, X Company's profits would fall by
Answer #1

Solution 6:

Variable cost of goods sold per unit = ($614,390 - $140,170) / 65500 = $7.24 per unit

Variable period cost per unit = ($176,850 - $81,875) / 65500 = $1.45 per unit

Variable period cost per unit of special order = $1.45 - ($19*4%) + ($12*4%) = $1.17 per unit

Profit on special order = Contribution margin per unit of special order * Nos of units sold

= ($12 - $7.24 - $1.17) * 4260 = $15,293.40

Solution 7:

Decrease in Variable cost of goods sold per unit = $0.80

Decrease in variable period cost per unit = $0.48

Effect of these changes on special order profit = ($0.48 + $0.80) * 4260 = $5,452.80 Increase in special order profit

Solution 8:

Contribution margin per unit on regular orders = $19 - $7.24 - $1.45 = $10.31 per unit

If regular sales fall by 700 units then company's profit will fall by = 700 * $10.31 = $7,217